OpenStudy (anonymous):

I am helping my daughter with her math. We are working on solving inequalities. our problem is like this, 2x+1>3; 1 I have a basic idea, but not sure how to show her how to show work.

OpenStudy (radar):

\[2x+1>3\]

OpenStudy (radar):

subtract 1 from both sides as if it was a regular equality equation. 2x>3-1 or 2x>2 divide both sides by 2 giving the answer x>1 if you divide by or multiply by a 1 change the direction of the inequality. did not have to in this problem, always check your results

OpenStudy (radar):

I meant to say a minus 1!!

OpenStudy (anonymous):

give me min to review

OpenStudy (radar):

when multiplying both sides by a negative change the direction of the inequality sign. Hope this did not confuse you

OpenStudy (anonymous):

The directions say to decide whether the given number is a solution of the inequality

OpenStudy (anonymous):

im still reviewing it. Im slow to understand math.

OpenStudy (radar):

No the given number shown as a 1 is not larger it actually equals 3

OpenStudy (anonymous):

but the x = 1?

OpenStudy (radar):

No look at my second post, x>1 not equal to one

OpenStudy (anonymous):

I dont see where you divided by 2. I see the rules but am confused

OpenStudy (anonymous):

so if the problem is 2x-1>3 you do it as if it were an equal sign so it be 2x-1=3 so you add one to the right side which will be 2x=4 so you divide by 2 which is x=2 but since it was a (>) its x>2 and since you didn't divide by a negative you don't chance the sign. i hope it helps

OpenStudy (anonymous):

the problem said +1, but you subtract it?

OpenStudy (anonymous):

I am confused, I dont see any negative numbers.

OpenStudy (radar):

I divided both sides by 2 where it was 2x>2 2x/2 was on one side and 2/2 was on the other giving x>1

OpenStudy (anonymous):

my bad so if its 2x+1>3 do it like an equal sign. 2x+1=3 so you subtract one to the right side which will make it 2x=2 so to leave the x by itself you divide by two. So it be x=1 but since you had(>) it be x>1

OpenStudy (radar):

Remember how you try to isolate the x on one side and have constants on the opposite side? If you subtract from one side you must subtract from the other side.

OpenStudy (radar):

I subtracted the 1 to isolate the unknown on one side.

OpenStudy (anonymous):

2x+1>3;1 is the same as 2x-3=1

OpenStudy (anonymous):

are you just dropping the +1

OpenStudy (radar):

No, look carefully 2x+1>3 is the same as x>1

OpenStudy (anonymous):

are you replacing the + with the >

OpenStudy (radar):

No I am doing a series of steps that will solve the problem. Step 1 was to subtract 1 from both sides giving 2x>3-1 or 2x>2 You do see that on the left side 2x+1-1 = 2x

OpenStudy (anonymous):

ok, thanks,

OpenStudy (anonymous):

2x+1>3 -1 -1 2x>2 /2 /2 x>1 Thats the answer

OpenStudy (radar):

Good luck teaching your daughter

OpenStudy (anonymous):

thanks, i need as much help as her LOL